OpenStudy (anonymous):

How many solutions does the following system have: y=12? Why?

OpenStudy (anonymous):

where is the system?

OpenStudy (anonymous):

Trying to solve a system of linear equations using substitution.

OpenStudy (anonymous):

yes but where is the system? the equations?

OpenStudy (anonymous):

Sorry, I meant to say the RESULT is y=12 for a system of unknown linear equations using substitution. How many solutions does the system have?

OpenStudy (unklerhaukus):

y=12 is one solution if there are no other solutions then there is only one solution

OpenStudy (anonymous):

two variable equations? im not sure i understand the question

OpenStudy (anonymous):

right unklerhaukus is right

OpenStudy (anonymous):

i just got confused for a sec

OpenStudy (anonymous):

Thanks to both of you. That makes sense.

OpenStudy (anonymous):

If the result of a system of linear equations is 8=8, why are the number of solutions infinite?

OpenStudy (anonymous):

because you know that the variables canceled themselves out such as y = 2x 2x + 8 = y + 8 8 = 8 you can put in any number for y or x (but not both) that you want and it will always be true, you cant do both because y = 2x therefore you cant do y = 10 and x = 8

OpenStudy (anonymous):

Great explanation mridrik. Thanks

OpenStudy (anonymous):

welcome